# SelectLLM: Can LLMs Select Important Instructions to Annotate?

### 29 Jan 2024 v1

@@ 4.1 Setups
Datasets and metrics. We use labeled datasets without using their responses/labels to test our hypothesis. Further, we utilize one human-generated dataset and one machine-generated dataset. For the former, we use (Conover et al., 2023) which is a combined effort of several Databricks employees, and as for the machine-generated dataset, we use Cleaned Alpaca which is based on (Taori et al., 2023) but with responses generated by GPT-4 and cleaned up to fix any errors in the prompts of the original dataset. When it comes to evaluating the performance, we analyze the similarity between the inferred text and the actual text using the Rouge L score. We also measure the cosine similarity between the inferred text and the actual text. Additionally, we perform a GPT-based analysis to gauge the effectiveness of the generated inferences.
@@
Baselines. We consider several baselines for comparison with our algorithm as follows: (1) Random: selecting instances from the unlabeled dataset purely at random. (2) Length: Considers the length of input instruction, focusing on both longer and shorter ones to evaluate their impact (Lengthlong and Lengthshort). (3) Cluster-Based Selection (CBS) (Chen et al., 2023a): transforming instructions into embedding space, clustering them with HDBSCAN, and selecting samples using the K-Center-Greedy algorithm. We consider two different embedding spaces with Sentence-bert (Reimers and Gurevych, 2019) and Instructor (Su et al., 2023), and denote them as CBSsbert and CBSinstr, respectively. (4) Perplexity (Marion et al., 2023): selecting samples based on low per-token perplexity, indicating high model certainty and fluency. (5) Diversity (Wang et al., 2022a): for each instruction in the dataset, Rouge score is computed against a randomly selected subset comprising $n$ samples ($n\ll M$). Then, we select $k$ samples that exhibit the minimum Rouge scores. (6) Open-Endedness (OpenEnd) (Li et al., 2023): generating three inferences per prompt, counting unique bigrams, and selecting samples with the greatest variety of bigrams. (7) Coreset (Sener and Savarese, 2018): similar to CBS, transforming instructions into embedding space with Sentence-bert, then selecting samples with K-Center-Greedy algorithm (Sener and Savarese, 2018).
@@
Table 2: Experimental results on Cleaned Alpaca (Conover et al., 2023). Rouge (F1) and Cosine similarity of generated responses from fine-tuned LLaMA-2 by different numbers of examples are compared. The best and second best scores are highlighted in bold and underline, respectively.
@@
Rouge (F1) Cosine Similarity Avg Across Sizes Methods $1k$ $2k$ $3k$ $1k$ $2k$ $3k$ Avg Rouge Avg Cosine Lengthshort 0.219 0.263 0.261 0.519 0.632 0.625 0.248 0.592 Perplexity 0.264 0.278 0.272 0.628 0.646 0.640 0.271 0.638 CBSsbert 0.254 0.264 0.288 0.598 0.618 0.665 0.269 0.627 Lengthlong 0.228 0.266 0.297 0.550 0.618 0.683 0.264 0.617 CBSinstr 0.257 0.280 0.292 0.610 0.655 0.673 0.276 0.646 Random 0.281 0.281 0.271 0.653 0.662 0.656 0.278 0.657 Diversity 0.268 0.286 0.297 0.650 0.673 0.690 0.284 0.671 OpenEnd 0.250 0.247 0.276 0.616 0.601 0.645 0.258 0.621 Coreset 0.277 0.267 0.299 0.650 0.635 0.695 0.281 0.660 Ours 0.276 0.277 0.301 0.661 0.667 0.707 0.285 0.678
Implementation details. For our experiment, we utilize the Dolly and Cleaned Alpaca datasets for both the training and evaluation phases. Since there is no explicit train and test split for these datasets, we reserve 1k samples from each dataset to form our test sets. This allocation leaves us with 51k samples in the Cleaned Alpaca dataset and 14k samples in the Dolly dataset for training purposes. Then, we apply one of our sampling algorithms to each training dataset to select a subset of data, varying the subset size between 1k to 3k samples for training, with an 80:20 training and validation split. We fine-tune LLaMA-2 (7B) model (Touvron et al., 2023) by employing QLoRa (Dettmers et al., 2023), a model optimization technique, to reduce the memory requirements during the fine-tuning process. For the experiments, we use three different random seeds and then compute the average of the evaluation scores from these three models to derive a final score for each method. We run a total of 20 epochs with a batch size of 6. We use the Paged optimizer and set the gradient accumulation steps at 2. To avoid overfitting and select the best model, we integrate an Early Stopping Callback with a patience of 3 epochs and a threshold of 0.01. Also, for selecting instructions with SelectLLM, we commonly use ChatGPT (gpt-3.5-turbo-0613).
@@ 4.2 Main Results
In this section, we present our main experimental results. Our comprehensive evaluation, incorporating both Rouge (F1) scores and Cosine Similarity metrics, provides a detailed insight into the performance variations across different sample sizes (1k, 2k, and 3k), as well as the average performance across these sizes. The results for Dolly are detailed in Table 1 and for Alpaca in Table 2, respectively. To be specific, our analysis leads to several nuanced observations:
@@
Consistent effectiveness across datasets. SelectLLM exhibits unparalleled consistency in both human and synthetic datasets. This uniformity across sample sizes sets it apart from other baselines and demonstrates its broad applicability. In contrast, other methods like Coreset, Diversity, and LengthBig show fluctuating performances depending on the dataset and sample size. For example, Coreset varies notably with sample size, while Diversity and LengthBig excel in the Cleaned Alpaca dataset but falter in the Dolly dataset. OpenEndedness, interestingly, performs better in Dolly but shows decreased effectiveness in Cleaned Alpaca. These variations in other methods further highlight the robust and adaptable nature of SelectLLM.
Table 3: Cross-dataset generalization. Rouge (F1) / Cosine similarity of generated responses from fine-tuned LLaMA-2 by selecting 3k examples on counterpart (e.g., train: Dolly $\rightarrow$ eval: Cleaned Alpaca) with different methods are compared. The best scores are highlighted in bold, and columns are the trained dataset.
@@
Cross-dataset generalization. We conducted an analysis to determine how well models trained on dolly samples using various sampling techniques generalize to cleaned alpaca data, and vice versa. Our findings are presented in Table 3 for both datasets. We made an intriguing observation with regards to models trained on dolly data - the model trained using our approach remarkably performed better than all the baselines by 10% on the cleaned alpaca test set. On the Cleaned Alpaca dataset, Coreset shows comparable performance to SelectLLM in terms of the Cosine Similarity and slightly better performance in rouge scores compared to SelectLLM. Further, the Cleaned Alpaca dataset appears to be a better dataset for cross-evaluation generalization when observing the performance of all baselines trained on it.
Evaluation with GPT-4. We evaluate the quality of generated responses between fine-tuned LLaMA-2 with SelectLLM and other methods with a 1k size dataset. Specifically, we randomly selected 1k size instructions, inputs from the original dataset, and generated responses from finetuned models. We ask GPT-4 to choose a better response to a given instruction, similar to (Liu et al., 2023). As shown in Table 4, when comparing SelectLLM to Random, SelectLLM won in 52% of the cases. We present the used prompt in Figure 6.
@@
Table 5: Ablation study. Rouge (F1) and cosine similarity of generated responses from fine-tuned LLaMA-2 by selecting 1k examples on Dolly with different methods are compared. The best scores are highlighted in bold.
@@ 4.3.1 Ablation study
To show the effectiveness of each component of SelectLLM, we ablate against different combinations of local selection prompting methods (Sec 3.2) and global division algorithms (Sec 3.3) on the 1k samples with Dolly.
Table 5 shows the results.
@@ 4.3.2 Qualitative results
Comparison of outputs from fine-tuned LLMs. Here, we compare the generated responses from fine-tuned LLMs using SelectLLM against two representative baselines, Random and Coreset. This comparison focuses on their adherence to and interpretation of instructions in both closed and open QA formats. The primary example of the analysis is detailed in Table 6, while additional examples are in the appendix (see Table 9).
In the example in Table 6, SelectLLM demonstrates superior precision and adherence to the instruction compared to the other methods. While the Coreset and Random methods provide correct responses, they tend to replicate the input text without much specificity. In contrast, SelectLLM generates a response that is not only precise and direct but also effectively extracts and succinctly presents the relevant information.
The additional examples, presented in Table 9, further underscore SelectLLM’s nuanced understanding and response capability. In an open-ended task, for instance, Random and Coreset deliver basic, concise responses, but lack the personalized and empathetic touches seen in SelectLLM’s reply. SelectLLM infuses its response with a sense of passion and commitment, demonstrating a deeper comprehension of the instruction’s intent. This sophistication in response is indicative of the advanced capabilities necessary in instruction-tuned models. Another closed QA example asking to differentiate between two terms shows all three models arriving at the correct answer, yet only SelectLLM clearly distinguishes between the two terms by providing detailed definitions, highlighting its capacity to deliver more informative and contextually rich responses.
@@ 5 Conclusion
In this work, we present SelectLLM, a method to utilize LLMs for selecting an effective subset among unlabelled instructions. Through the various experiments based on two popular benchmarks, we validate the effectiveness of SelectLLM compared to the previous selection methods. This demonstrates its effectiveness in improving the efficiency of instruction-tuning for language models.
@@
The findings from our research not only establish the proficiency of LLMs in autonomously selecting high-quality data for training but also open new paths for future investigation in this domain. The successful application of LLMs in data-constrained environments is demonstrated by the exceptional ability of SelectLLM. This study, therefore, marks a significant stride in the field of Instruction-tuning for LLMs, paving the way for more efficient and effective training methodologies and expanding the scope of autonomous capabilities of LLMs.
Acknowledgements
The work has been supported by funding from Cisco. We would like to thank Cisco team and Minnesota NLP group members for providing us with valuable feedback and comments on the initial draft.
@@ A.2 Baselines
2) Coreset-Based Selection (Chen et al., 2023a): Uses clustering for sample selection. Involves transforming instructions into vector representations, clustering them with HDBSCAN, and selecting samples using the K-Center-Greedy algorithm which focuses on cluster centroids.
@@
5) Open-Endedness (Li et al., 2023): Determines prompt openness by generating three inferences per prompt, counting unique bigrams, and selecting samples with the greatest variety of bigrams. This process follows the open-endedness criteria defined for samples with a broader range of chain of thought reasonings in (Li et al., 2023).
[…]
7) Coreset (Sener and Savarese, 2018): Similar to CBS, transforming instructions into embedding space with Sentence-bert, then selecting samples with K-Center-Greedy algorithm.
8) Length: Considers the length of prompts, focusing on both longer and shorter ones to evaluate their impact.
@@ Appendix C Verifying Capability of LLMs for Selecting Unlabelled Instruction
Table 9: More examples of generated sentences.
Hi,
I hope you are having a wonderful day. My name is _, and I just applied for the Program Staff position at the Services for Students with Disabilities to empower students who have disabilities to reach their academic dreams. My background working in higher education and volunteering with kids on the Autism spectrum combine to bring an empathetic and impactful perspective.
Please let me know if there is anything else I can do at this time to be considered for this position.
With gratitude,
_
Random Coreset Ours Dear Head of Academic Coaching,
Thank you for your time.
Sincerely,
Katie

### 20 Feb 2024 v2

@@ 4.1 Setups
Datasets and metrics. We use labeled datasets without using their responses to test our hypothesis. Further, we utilize one human-generated dataset and one machine-generated dataset. For the former, we use (Conover et al., 2023) which is a combined effort of several Databricks employees, and as for the latter, we use Cleaned Alpaca which is based on (Taori et al., 2023) but is cleaned up to fix any errors in the input prompts of the original dataset and has responses generated by GPT-4. For evaluating the performance, we assess the similarity between inferred and actual texts using Rouge scores and cosine similarity. Additionally, we perform a GPT-based analysis to gauge the effectiveness of the generated inferences.
@@
Baselines. We consider several baselines for comparison with our algorithm as follows: (1) Random: selecting instances from the unlabeled dataset purely randomly. (2) Length: Considers the length of input instruction, focusing on both longer and shorter ones to evaluate their impact (Lengthlong and Lengthshort). (3) Cluster-Based Selection (CBS) (Chen et al., 2023a): transforming instructions into embedding space, clustering them with HDBSCAN (Campello et al., 2013), and selecting samples using the K-Center-Greedy algorithm. We consider two different embedding spaces with Sentence-BERT (Reimers and Gurevych, 2019a) and InstructOR (Su et al., 2023), and denote them as CBSsbert and CBSinstr, respectively. (4) Perplexity (Marion et al., 2023): selecting samples based on low per-token perplexity, indicating high model certainty and fluency. (5) Diversity (Wang et al., 2022a): for each instruction in the dataset, Rouge score is computed against a randomly selected subset comprising $n$ samples ($n\ll M$). Then, we select $k$ samples that exhibit the minimum Rouge scores. (6) Open-Endedness (OpenEnd) (Li et al., 2023): generating three inferences per prompt, counting unique bigrams, and selecting samples with the greatest variety of bigrams. (7) Coreset (Sener and Savarese, 2018): Similar to CBS, transforming instructions into embedding space with Sentence-bert, then selecting samples with K-Center-Greedy algorithm (Sener and Savarese, 2018).
@@
Table 2: Experimental results on Cleaned Alpaca (Conover et al., 2023). Rouge-L (F1) and Cosine similarity of generated responses from fine-tuned LLaMA-2 by different numbers of examples are compared. The best and second best scores are highlighted in bold and underline, respectively.
@@
Rouge-L (F1) Cosine Similarity Avg Across Sizes Methods $1k$ $2k$ $3k$ $1k$ $2k$ $3k$ Avg Rouge Avg Cosine Lengthshort 0.219 0.263 0.261 0.519 0.632 0.625 0.248 0.592 Perplexity 0.264 0.278 0.272 0.628 0.646 0.640 0.271 0.638 CBSsbert 0.254 0.264 0.288 0.598 0.618 0.665 0.269 0.627 Lengthlong 0.228 0.266 0.297 0.550 0.618 0.683 0.264 0.617 CBSinstr 0.257 0.280 0.292 0.610 0.655 0.673 0.276 0.646 Random 0.281 0.281 0.271 0.653 0.662 0.656 0.278 0.657 Diversity 0.268 0.286 0.297 0.650 0.673 0.690 0.284 0.671 OpenEnd 0.250 0.247 0.276 0.616 0.601 0.645 0.258 0.621 Coreset 0.277 0.267 0.299 0.650 0.635 0.695 0.281 0.660 Ours 0.276 0.277 0.301 0.661 0.667 0.707 0.285 0.678
Implementation Details. We utilize the Dolly and Cleaned Alpaca datasets for training and evaluation. Due to the absence of predefined train-test splits in these datasets, we allocated 1k samples from each for testing, leaving 14k and 51k samples in the Dolly and Cleaned Alpaca datasets, respectively, for training. We employed sampling algorithms to select subsets of 1k to 3k samples from each train set. The LLaMA-2 (7B) model was fine-tuned using QLoRa (Touvron et al., 2023) to optimize memory usage during training. For instruction selection with SelectLLM, ChatGPT (gpt-3.5-turbo-0613) was utilized. To ensure robustness in our results, we conducted experiments using three different random seeds. The models’ evaluation scores were averaged to derive a final score for each method. Further details regarding the training process are provided in the Appendix.
@@ 4.2 Main Results
In this section, we present our main experimental results. Our comprehensive evaluation, incorporating both Rouge scores and Cosine Similarity metrics, provides a detailed insight into the performance variations across different sample sizes (1k, 2k, and 3k), as well as the average performance across these sizes. The results for Dolly are detailed in Table 1 and for Alpaca in Table 2, respectively. To be specific, our analysis leads to several nuanced observations:
@@
Consistent effectiveness across datasets. SelectLLM exhibits unparalleled consistency in both human and synthetic datasets. This uniformity across sample sizes sets it apart from other baselines and demonstrates its broad applicability. In contrast, other methods like Coreset, Diversity, and Lengthlong show fluctuating performances depending on the dataset and sample size. For example, Coreset varies notably with sample size, while Diversity and Lengthlong excel in the Cleaned Alpaca dataset but falter in the Dolly dataset. OpenEndedness performs better in Dolly but shows decreased effectiveness in Cleaned Alpaca. This further highlights the robust and adaptable nature of SelectLLM.
Cross-dataset generalization. We analyze how well models trained on dolly samples using various sampling techniques generalize to cleaned alpaca data, and vice versa. Our results are presented in Table 3 for both datasets. We made an intriguing observation with regards to models trained on dolly data - the model trained using our approach remarkably performed better than all the baselines by 10% on the cleaned alpaca test set. On the Cleaned Alpaca dataset, Coreset shows comparable performance to SelectLLM in terms of Cosine Similarity and slightly better performance in Rouge scores. Further, Cleaned Alpaca appears to be a better dataset for cross-evaluation generalization when observing the performance of all baselines trained on it.
Table 3: Cross-dataset generalization for 3k sample size. Rouge-L (F1) / Cosine similarity of generated responses from LLaMA-2. The best scores are highlighted in bold, and column names indicate the dataset trained on.
@@
Evaluation with GPT-4. We evaluate the quality of generated responses between LLaMA-2 fine-tuned with SelectLLM and other methods on a 1k size instruction dataset, randomly sampled from the Dolly dataset. We ask GPT-4 to choose the better response to a given instruction, similar to (Liu et al., 2023). As shown in Table 4, SelectLLM wins in 52% of the cases when compared to Random sampling, and 44% cases when compared to Coreset sampling, further showcasing better inference quality of SelectLLM. The prompt for the evaluation is provided in the appendix (Figure 6).
@@
Table 5: Ablation study. Rouge and cosine similarity of generated responses from fine-tuned LLaMA-2 by selecting 1k examples on Dolly with different methods are compared. The best scores are highlighted in bold.
@@
Hi,
I hope you are having a wonderful day. My name is _, and I just applied for the Program Staff position at the Services for Students with Disabilities to empower students who have disabilities to reach their academic dreams. My background working in higher education and volunteering with kids on the Autism spectrum combine to bring an empathetic and impactful perspective. Please let me know if there is anything else I can do at this time to be considered for this position. With gratitude,
Random Coreset Ours Dear Head of Academic Coaching,
Thank you for your time.
Sincerely,
Katie
@@ 4.3.1 Ablation study
To show the effectiveness of each component of SelectLLM, we ablate against different combinations of local selection prompting methods (Sec 3.2) and global division algorithms (Sec 3.3) on 1k samples from Dolly. Table 5 shows the results.
@@ 4.3.2 Qualitative results
Comparison of outputs from fine-tuned LLMs. We evaluate responses from LLMs fine-tuned with SelectLLM against Random and Coreset baselines, focusing on their handling of closed and open QA formats. Key examples are in Table 6, with more in the appendix (Table 9).
In the first example in Table 6, SelectLLM demonstrates superior precision and adherence to the instruction compared to the other methods. While the Coreset and Random methods provide correct responses, they tend to replicate the input text without much specificity. In contrast, SelectLLM generates a response that is not only precise and direct but also effectively extracts and succinctly presents the relevant information.
The second example, based on an Open-ended task further underscores SelectLLM’s nuanced understanding and response capability. While Random and Coreset give basic, concise answers, SelectLLM adds personalized, empathetic elements (highlighted in blue), showing deeper instruction comprehension. This reflects the advanced response capabilities needed in instruction-tuned models.
@@ 5 Conclusion
We introduce SelectLLM, a new approach that uses LLMs to choose an efficient subset from a set of unlabelled instructions. Our experiments on two popular benchmarks show that SelectLLM is more effective than previous selection methods. This demonstrates how LLMs can improve the efficiency of instruction tuning for language models.
@@
The findings from our research not only establish the proficiency of LLMs in autonomously selecting high-quality data for training but also open new paths for future investigation in this domain. The successful application of LLMs in data-constrained environments is demonstrated by the exceptional ability of SelectLLM. This study, therefore, marks a significant stride in the field of instruction tuning for LLMs, paving the way for more efficient and effective training methodologies and expanding the scope of autonomous capabilities of LLMs. In terms of ethical implications, the potential for any risk is limited to the application of LLMs in our framework, and the general risks associated with them such as LLMs showing bias in selecting certain instructions according to what it believes to be an impactful instruction. Further, bias can also be introduced based on how the prompt is designed by the user, when querying the LLMs in our framework.
@@ A.2 Baselines
[…]
@@
5) Open-Endedness (Li et al., 2023): Determines prompt open-endedness by generating three inferences per prompt, counting unique bigrams between the generated inferences, and selecting samples with the greatest variety of bigrams. This process follows the open-endedness criteria defined for samples with a broader range of chain of thought reasonings in (Li et al., 2023).
6) Coreset (Sener and Savarese, 2018): Similar to CBS, this method involves transforming data points into their embedding space with Sentence-BERT, then iterates a process to get subsets that maximize the coverage and diversity within each subset with a predetermined subset size. The algorithm selects samples by prioritizing those that maximize the distance to the nearest point already included in the subset, ensuring that the selected samples are diverse within each subset.
7) Length: Considers the length of prompts (Instruction + Input), focusing on both longer and shorter ones to evaluate their impact.
@@ A.3 SelectLLM
A.4 Implementation details
For our experiment, we utilize the Dolly and Cleaned Alpaca datasets for both the training and evaluation phases. As there were no explicit train and test split for these datasets, we randomly sampled 1k samples from each dataset to form our test sets. This allocation leaves us with 51k samples in the Cleaned Alpaca dataset and 14k samples in the Dolly dataset for training purposes. Then, we apply one of our sampling algorithms to each training dataset to select a subset of data, varying the subset size between 1k to 3k samples for training, with an 80:20 training and validation split. We fine-tune LLaMA-2 (7B) model (Touvron et al., 2023) by employing QLoRa (Dettmers et al., 2023), a model optimization technique, to reduce the memory requirements during the fine-tuning and inference processes. For the experiments, we use three different random seeds and then compute the average of the evaluation scores from these three models to derive a final score for each method. We run a total of 20 epochs with a batch size of 6. We use the Paged optimizer and set the gradient accumulation steps at 2. To avoid overfitting and select the best model, we integrate an Early Stopping Callback with a patience of 3 epochs and a threshold of 0.01. Also, for selecting instructions with SelectLLM, we commonly use ChatGPT (gpt-3.5-turbo-0613).
@@ Appendix C Verifying Capability of LLMs for Selecting Unlabelled Instruction
Table 9: Another instance of a closed QA example asking to differentiate between two terms. All three models arrive at the correct answer, however only SELECTLLM clearly distinguishes between the two terms by providing detailed definitions, highlighting its capacity to deliver more informative and contextually rich explanations
@@ Appendix D More Analyses with Chosen Instructions
Table 10: Experimental results on Dolly (Conover et al., 2023). Rouge-L (F1) and Cosine similarity of generated responses from fine-tuned LLaMA-2 models on the entire dolly dataset are compared with a model trained on only 3k samples chosen by SelectLLM.
Methods Rouge-L (F1) Cosine Similarity Dolly-15k 0.292 0.689 SelectLLM-3k 0.289 0.686
Table 11: Experimental results on Cleaned Alpaca (Conover et al., 2023). Rouge-L (F1) and Cosine similarity of generated responses from fine-tuned LLaMA-2 models on the entire Cleaned Alpaca dataset is compared with a model trained on only 3k samples chosen by SelectLLM.
Methods Rouge-L (F1) Cosine Similarity Dolly-15k 0.292 0.689 SelectLLM-3k 0.289 0.686
@@
Appendix E Comparison of Selected Sample with the Entire Dataset
We compared the SelectLLM-3k sampled model with models trained on full Dolly and Cleaned Alpaca datasets. As detailed in 10 and 11, SelectLLM-3k nearly equals the full Dolly dataset model in performance, and achieves 85% of the full dataset’s performance on the Cleaned Alpaca dataset, as per the Rouge score. This underscores SelectLLM’s efficiency in filtering out noise from the entire dataset, especially in the case of Dolly. While our results are akin to the data filtering approach in (Chen et al., 2023b), which also samples 3k instructions from the Dolly dataset, there are notable differences. Their method outperforms the full dataset, possibly due to their inclusion of sample outputs in the filtering process, an aspect not considered in our approach.